If you begin with 995 grams of ch4 and 2510 grams of h2o. How many grams of CO could you make from this reaction?
1
Expert's answer
2020-03-24T14:02:34-0400

CH4 + H2O "\to" CO + 3H2

"n(CH_4)=m(CH_4)\/M(CH_4)"

"n(CH_4)=995g\/(16g\/mole)=62.19 mole"

"n(H_2O)=m(H_2O)\/M(H_2O)"

"n(H_2O)=2510g\/(18g\/mole)=139.44mole"

As we can see methane is in short supply.

According to the equation n(CH4)= n(CO)

"m(CO)=62.19mole*28g\/mole=1741.32g"

The answer is 1741.32 grams.